Please show me how to solve the inequality stating the solution set using intervla notation and graph the solution set. Any help and explanation will be appreciated.
4x^2 - 8x ≥ 0

Answer by venugopalramana(3286)

4x^2-8x > = 0
4X(X-2) > =0
EITHER X=0 OR X=2...FOR 4x^2-8x = 0..SINCE A PRODUCT IS ZERO IF EITHER OF THE 2 FACTORS IS ZERO.
FOR 4x^2-8x > 0....EITHER BOTH FACTORS SHOULD BE POSITIVE OR NEGATIVE.
X > 0 AND X > 2....OR....X < 0 AND X < 2...THAT IS.......
X > 2...............OR.......X < 0